OpenStudy (cookiemonster18):

Can you help me check 2 questions? suppose y varies directly with X write a direct variation equation that relates X to Y Y=7 1/2 when X=3 A.y=3/2x B.y=2/3x C.y=5/2x D.y=2/5x (I think it's C am I correct?)

OpenStudy (cookiemonster18):

@rebelbluebeetle can you help me

OpenStudy (nikato):

C is correct since y vairies directly with x means y=kx

OpenStudy (nikato):

plug in the y and x and solve for k 7 1/2= k(3) 5/2 =k so your equation would be y=5/2 x

OpenStudy (cookiemonster18):

@nikato can you help with the 2nd one?